# AutoRev: Automatic Peer Review System for Academic Research Papers

### v1

@@
Abstract
Generating a review for an academic research paper is a complex task that requires a deep understanding of the document’s content and the interdependencies between its sections. It demands not only insight into technical details but also an appreciation of the paper’s overall coherence and structure. Recent methods have predominantly focused on fine-tuning large language models (LLMs) to address this challenge. However, they often overlook the computational and performance limitations imposed by long input token lengths. To address this, we introduce AutoRev, an Automatic Peer Review System for Academic Research Papers. Our novel framework represents an academic document as a graph, enabling the extraction of the most critical passages that contribute significantly to the review. This graph-based approach demonstrates effectiveness for review generation and is potentially adaptable to various downstream tasks, such as question answering, summarization, and document representation. When applied to review generation, our method outperforms SOTA baselines by an average of 58.72% across all evaluation metrics. We hope that our work will stimulate further research in applying graph-based extraction techniques to other downstream tasks in NLP. We plan to make our code public upon acceptance.
AutoRev: Automatic Peer Review System for Academic Research Papers
Maitreya Prafulla Chitale1thanks: Equal contribution.  Ketaki Mangesh Shetye [^1] footnotemark: 1 Harshit Gupta1Manav Chaudhary1Vasudeva Varma1 1IIIT Hyderabad {maitreya.chitale, ketaki.shetye, harshit.g, manav.chaudhary}@research.iiit.ac.in vv@iiit.ac.in
1 Introduction
Figure 1: Example graph constructed out of a paper consisting of 2 Headings, 2 Sub-Headings, 4 Passages, and 6 Sentences.
@@ 1 Introduction
Recent research has increasingly focused on developing automatic systems for peer review, including tools that assist reviewers, assess peer-review processes, and directly generate reviews for academic papers Yu et al. (2024); Aitymbetov and Zorbas (2025); Church et al. (2025). Concurrently, there has been growing interest in examining the strengths and limitations of such systems, particularly the use of large language models (LLMs) for automated review generation Ye et al. (2024). Motivated by both of these trends, we propose AutoRev, a system designed to generate peer reviews from a guidance perspective, offering constructive feedback to support researchers and reviewers.
@@
With the rise of Graph Neural Networks, the use of graphs as extractors or as guides for facilitating downstream tasks has recently gained traction in the NLP community Chitale et al. (2025); Gao et al. (2025); Agarwal et al. (2025). Inspired by this, we propose to use graphs in the domain of academic paper analysis, specifically, generating reviews that closely replicate those produced by existing peer review systems in conferences. AutoRev can generate high-quality reviews without relying on any prior human evaluations of the paper. Furthermore, by leveraging the graph’s capability as an extractor of passages, we significantly reduce the input token length to the LLM. Reducing the token length of the input has been shown to enhance LLM performance Hsieh et al. (2024).
To summarize our main contributions:
@@
  • We propose a system for generating concise, structured reviews (Summary, Strengths, Weaknesses, and Questions) to assist authors and reviewers.
  • To our knowledge, this is the first approach using graphs in this domain.
@@
  • Our approach improves over baselines by 58.72% on average across all metrics, without relying on human-written reviews and staying efficient.
@@
It is important to highlight that this paper does not seek to determine the final reviews of a submission. Rather, we present our framework as a tool to provide timely feedback to researchers and serve as a valuable guide for reviewers.
2 Related Works
@@
A significant amount of research has been conducted on aiding peer reviews for academic papers. Some studies explore the involvement of AI in the selection process Aitymbetov and Zorbas (2025), while some studies work on generating meta-reviews Kumar et al. (2023). Sufficient discussion and avenues have been explored where AI/NLP techniques can contribute towards aiding automatic peer reviews Kuznetsov et al. (2024). Though, to the best of our knowledge, significant research has not been done in the field of generating a whole review for an academic paper.
A recent study by Faizullah et al. (2024) explores a DPR-based approach for generating suggestive limitations. It highlights the necessity of using DPR to address length constraints and extract only the most relevant passages from a paper. This study served as an initial motivation for incorporating DPR into our approach.
Another notable work by Yu et al. (2024) introduces an automated paper reviewing framework. The work also includes a standardized dataset of GPT-4 OpenAI et al. (2024) consolidated reviews collected from various academic venues. To enhance review generation, the authors propose a self-correction strategy during LLM fine-tuning to ensure a higher quality of the generated reviews. We adopt this work as our baseline and aim to develop a more computationally efficient approach than direct LLM fine-tuning while achieving higher-quality reviews. The dataset publicised by the authors of this study is being directly used by us in our work.
Ye et al. (2024) highlight key issues in LLM-generated reviews, such as giving overly high ratings in some cases and being easily influenced by authors to generate only minor limitations. Considering the contributions of this work, we reinforce the need for our approach, not as a tool for final review decisions but as a means to provide constructive feedback to researchers and reviewers.
3 Dataset
@@
The dataset used in our study is borrowed from Yu et al. (2024), comprising papers submitted to various conferences along with multiple peer reviews per paper. Among these, only the datasets for ICLR 2024 and NeurIPS 2023 are publicly available. However, the number of GPT-4-consolidated reviews for NeurIPS 2023 is limited to just 235 samples. Hence, our analysis focuses only on the ICLR 2024 papers. As reported in Yu et al. (2024), the dataset statistics for the ICLR 2024 papers are summarized in Table 1.
@@
Table 1: Dataset statistics for the ICLR 2024 dataset, as reported in Yu et al. (2024).
@@
From the dataset, we utilize only the Summary, Strengths, Weaknesses, and Questions sections from each review, along with the full content of the corresponding paper. Other sections, such as Ratings are excluded, as our framework is designed to provide constructive feedback rather than produce a final decision. The selected sections offer qualitative assessments of the paper’s content and provide valuable insights to guide both authors and reviewers. Additionally, we use the GPT-4 OpenAI et al. (2024) consolidated reviews dataset, sourced from Yu et al. (2024), for finetuning LLMs to generate final structured reviews. This dataset covers 20% of randomly selected ICLR-24 papers (2036 papers), where the multiple peer reviews are consolidated using GPT-4 to generate integrated reviews in a unified format.
4 Methodology
In this section, we describe the implementation of our proposed framework. Section 4.1 describes the initial passage extraction using Dense Passage Retrieval (DPR) Karpukhin et al. (2020). These extracted passages are used as training data in Section 4.3, to train a Graph Neural Network (GNN) that learns to extract relevant passages from a research paper. This trained GNN enables our model to perform competitively on the test set without the need for grounding on human-generated reviews. Post that, we describe our process of fine-tuning LLMs to generate the final reviews, in Section 4.5. An overview of our methodology can be found in Figure 2
Figure 2: The system begins with a parsed document from which DPR extracts passages for GNN training. The GNN learns to identify key passages relevant to review generation. The trained GNN selects important passages without requiring human-generated reviews. These passages, combined with a structured prompt, are input to fine-tune an LLM. The fine-tuned LLM generates the final review efficiently, without external grounding.
4.1 DPR-based Passage Extraction
@@
This final set of passages represents the most relevant excerpts from the full paper that are semantically linked to the review content. These extracted passages serve as the gold-standard outputs for training the GNN, as described in Section 4.3
4.2 Graph Construction for Representing a Paper
The idea of graph construction employed in our work is inspired by Zhao et al. (2024). We borrow their idea of representing a document as a hierarchical graph but extend it beyond the sentence level. We propose to build a deeper graph, which contributes towards learning more intricate inter-dependencies of passages and sentences in a paper.
Specifically, we formalize the document structure as an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents nodes and $\mathcal{E}$ denotes edges encoding hierarchical and sequential relationships.
Node Definitions: Our graph construction is designed to be generalizable across diverse academic paper structures, accommodating arbitrary levels of nesting and heterogeneous section organization.
@@ 4.2 Graph Construction for Representing a Paper
  • Heading Nodes $\mathcal{H}={h_{1},h_{2},\dots,h_{m}}$: Nodes corresponding to top-level section headings.
  • Sub-heading Nodes $\mathcal{S}_{i}={s_{i1},s_{i2},\dots,s_{ik}}$: Nodes representing sub-headings under each heading $h_{i}$. This structure naturally extends to accommodate multiple levels of nesting, making it suitable for documents with deeper hierarchical organization.
@@
  • Sentence Nodes $\mathcal{T}_{jk}={t_{jk1},t_{jk2},\dots,t_{jkp}}$: Nodes for individual sentences within each passage $o_{jk}$, tokenized using NLTK Bird and Loper (2004).
Edge Definitions: Edges capture both hierarchical relationships and sequential flow within the document.
  • Root connections:
    $(v_{\text{paper}},h_{i})\in\mathcal{E}_{\text{hier}},\ \forall h_{i}\in\mathcal{H}$
  • Heading-subheading:
    $(h_{i},s{ij})\in\mathcal{E}_{\text{hier}},\ \forall s{ij}\in\mathcal{S}i$
@@
    $(s_{ij},o_{jk})\in\mathcal{E}_{\text{hier}}\text{ if }\mathcal{S}_{i}\neq\emptyset;\\ (h_{i},o_{jk})\in\mathcal{E}_{\text{hier}}\text{ otherwise}$
  • Passage-sentence:
    $(o_{jk},t_{jkq})\in\mathcal{E}_{\text{hier}},\ \forall t_{jkq}\in\mathcal{T}_{jk}$
  • Sequential edges (horizontal flow):
    • Passages: $(o_{jk},o_{j(k+1)})\in\mathcal{E}_{\text{seq}},\ \forall\ 1\leq k<n$
    • Sentences: $(t_{jkq},t_{jk(q+1)})\in\mathcal{E}_{\text{seq}},\ \forall\ 1\leq q<p$
@@
$\mathcal{E}=\mathcal{E}_{\text{hier}}\cup\mathcal{E}_{\text{seq}}$ (1)
@@
4.3 GNN Training for Passage Retrieval
To train a GNN on the graph constructed in Section 4.2, we take the data generated in Section 4.1, which maps each paper $p_{i}$ to a set of important passages, $\mathcal{O}_{i}=\{o_{1},o_{2},\dots,o_{n}\}$. Each passage is important for generating a review for the paper it is associated with. This mapping is used as a training set for training the GNN.
The node embeddings for sentence nodes are initialized with embeddings from a Sentence-Transformer Reimers and Gurevych (2019) model, namely all-mpnet-base-v2. The node embeddings for all other nodes are initialized with zero embeddings of dimension $768$, matching that of the Sentence-Transformer model.
Now, a Graph Attention Network Veličković et al. (2018) is used to train the graph to predict the important passages. After passing the initial representation of the graph nodes through three layers of GATConv [^3] https://pytorch-geometric.readthedocs.io/en/latest/generated/torch_geometric.nn.conv.GATConv.html (PyTorch Geometric’s implementation of Graph Attention Networks), we apply a non-linear classifier to the final passage node representations. The classification is defined as follows:
@@ 4.3 GNN Training for Passage Retrieval
We hypothesize that our choice of graph construction allows our model to capture the intricate inter-dependencies of sentences and passages across different sections within an academic paper, and effectively make a choice about which passage(s) is/are important. The dependencies, according to our hypothesis, will be learned from the sequential and hierarchical connections that we have introduced in our graph construction method. The passing of information from one node to another will be enhanced by the GATConv (Equations 2, 3, and 4) layers, which involves the concept of attention to selectively propagate messages in the Message Passing layer of a GNN.
$\mathbf{h}_{i}^{\prime}=\sigma\left(\sum_{j\in\mathcal{N}(i)}\alpha_{ij}\mathbf{W}\mathbf{h}_{j}\right)$ (2)
$e_{ij}=\text{LeakyReLU}\Bigl{(}\mathbf{a}^{T}\bigl{[}\mathbf{W}\mathbf{h}_{i}\parallel\mathbf{W}\mathbf{h}_{j}\bigr{]}\Bigr{)}$ (3)
$\alpha_{ij}=\frac{\exp\bigl{(}e_{ij}\bigr{)}}{\sum_{k\in\mathcal{N}(i)}\exp\bigl{(}e_{ik}\bigr{)}}$ (4)
@@
Figure 3: Excerpt of a test sample where passages highlighted in cyan colour refer to relevant passages identified by GNN only and yellow colour refer to relevant passages identified by both GNN and DPR
As illustrated in an example from our test set (Figure 3), the DPR-based approach fails to retrieve a key passage outlining the main contributions in the paper, whereas GNN successfully identifies it. Thus, leveraging sequentially and hierarchically connected graph nodes and GAT in GNN proves to be more effective in identifying important passages that help to enhance the review generation process. The GNN extracted passages, along with a structured prompt, serve as the input to LLM fine-tuning, as outlined in Section 4.5.
4.4 Grounding Requirements and Inference Efficiency
The advantage of training a GNN for this important classification task is twofold. First, unlike DPR-based approaches which require grounding at inference time via access to human-generated reviews or synthesized queries, our GNN-based method operates in a fully autonomous manner. Once trained, the GNN identifies salient passages directly from the graph-structured representation of the paper, eliminating the need for any external review or human guidance during inference. This distinction is crucial for real-world deployment, where human reviews may not be available or feasible to use.
Second, by selecting only the passages deemed most relevant, we significantly reduce the input length and associated computational costs for downstream LLM-based review generation. This not only leads to higher performance of LLMs, as outlined by Hsieh et al. (2024) but also leads to higher scores of the generated reviews (compared to our baselines), as reflected in our performance metrics (see Section 5). The statistics related to the decrease in number of passages and token length post-extraction by the GNN Model are reported in Table 2.

### v2

# AutoRev: Multi-Modal Graph Retrieval for Automated Peer-Review Generation

@@
Abstract.
Enhancing the quality and efficiency of academic publishing is critical for both authors and reviewers, as research papers are central to scholarly communication and a major source of high-quality content on the web. To support this goal, we propose AutoRev, an automatic peer-review system designed to provide actionable, high-quality feedback to both reviewers and authors. AutoRev leverages a novel Multi-Modal Retrieval-Augmented Generation (RAG) framework that combines textual and graphical representations of academic papers. By modelling documents as graphs, AutoRev effectively retrieves the most pertinent information, significantly reducing the input context length for LLMs and thereby enhancing their review generation capabilities. Experimental results show that AutoRev outperforms state-of-the-art baselines by up to 58.72% and demonstrates competitive performance in human evaluations against ground truth reviews. We envision AutoRev as a powerful tool to streamline the peer-review workflow, alleviating challenges and enabling scalable, high-quality scholarly publishing. By guiding both authors and reviewers, AutoRev has the potential to accelerate the dissemination of quality research on the web at a larger scale. Code will be released upon acceptance.
Graph Neural Networks (GNNs), Retrieval Augmented Generation (RAG), Large Language Models (LLMs), Automatic Peer Reviewing
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Information extractionccs: Computing methodologies Natural language generationccs: Information systems Language modelsccs: Information systems Document structure
1. Introduction
Figure 1. Example document graph with nodes for headings, sub-headings, passages, and sentences. This structure allows AutoRev to capture hierarchical and sequential dependencies across sections.:
@@ 1. Introduction
Recent research has increasingly focused on developing automatic systems for peer review, including tools that assist reviewers, assess peer-review processes, and directly generate reviews for academic papers (Yu et al., 2024; Aitymbetov and Zorbas, 2025; D’Arcy et al., 2024; Idahl and Ahmadi, 2025). Concurrently, there has been growing interest in examining the strengths and limitations of such systems, particularly the use of large language models (LLMs) for automated review generation (Ye et al., 2024). Motivated by both of these trends, we propose AutoRev, a system designed to generate peer reviews from a guidance perspective, offering constructive feedback to support researchers and reviewers.
@@
With the rise of Graph Neural Networks, the use of graphs as extractors or as guides for facilitating downstream tasks has recently gained traction in the NLP community (Chitale et al., 2025; Gao et al., 2025; Agarwal et al., 2025; Zhu et al., 2025). Inspired by this, we propose to use graphs in the domain of academic paper analysis, specifically, in generating reviews that closely replicate the structure and quality of those produced by existing peer review systems in conferences. Furthermore, by leveraging the graph’s capability as an extractor of passages, we significantly reduce the input token length to the LLM. Reducing the token length of the input is shown to enhance LLM performance (Hsieh et al., 2024). To summarize our main contributions:
  1. (1)
@@
    We propose a system for generating concise, structured reviews (Summary, Strengths, Weaknesses, and Questions) to assist authors and reviewers. To our knowledge, this is the first approach which uses graphs in this domain.
  2. (2)
@@
    Our approach improves over SoTA baselines by 58.72% on average across all metrics while being computationally efficient.
  3. (3)
    We conduct thorough human evaluations and other automatic evaluations to demonstrate the competitive performance of our method against ground truth reviews.
@@
It is important to highlight that this paper does not seek to determine the final reviews of a paper. Rather, we present our framework as a tool to provide timely feedback to researchers and serve as a valuable guide for reviewers.
We envision AutoRev as a step towards scalable, automated peer-review systems that can complement human judgment. With the continued surge in conference submissions across disciplines, the demand for such tools is increasingly pressing [^1] https://aaai.org/aaai-launches-ai-powered-peer-review-assessment-system/. By alleviating the bottlenecks of the traditional peer-review process, AutoRev can help accelerate the dissemination of high-quality research on the web, ensuring that credible scholarly knowledge remains accessible at scale.
2. Related Works
@@
A significant amount of research has been conducted on aiding peer reviews for academic papers. Some studies explore the involvement of AI in the reviewer selection process (Aitymbetov and Zorbas, 2025), while others focus on generating meta-reviews (Kumar et al., 2023). Prior work has also examined how AI/NLP techniques can contribute to various aspects of the peer-review pipeline (Kuznetsov et al., 2024).
Faizullah et al. (2024) explore a DPR-based approach for generating suggestive limitations, highlighting the necessity of retrieval methods to overcome length constraints. Building on this, we use DPR outputs as the training data for our Graph Neural Network (GNN), enabling it to learn to identify salient passages for review generation. Yu et al. (2024) introduce an automated paper reviewing framework (SEA) that standardizes reviews using GPT-4222https://arxiv.org/abs/2303.08774 consolidation and proposes a self-correction strategy during fine-tuning. Ye et al. (2024) further caution that LLM-generated reviews may overrate papers or produce superficial feedback, reinforcing our design goal of providing a structured and constructive review rather than final accept/reject decisions.
Among the most relevant recent systems, MARG (D’Arcy et al., 2024) employs a multi-agent framework, distributing different sections of a paper to specialized GPT-4 agents. This enables the model to handle long papers and produce more specific comments, but it comes at the cost of heavy computation and frequent inter-agent coordination errors. OpenReviewer (Idahl and Ahmadi, 2025) instead fully fine-tunes Llama-3.1-8B-Instruct on 79k reviews from ICLR and NeurIPS to generate structured reviews, adhering to the conference guidelines. This method of large-scale full fine-tuning is very costly.
We adopt SEA (Yu et al., 2024) as our primary baseline since it is closest in scope to our work, representing a recent state-of-the-art framework for automated peer review generation on the same ICLR dataset. Other systems such as MARG and OpenReviewer differ substantially in design. MARG uses multi-agent GPT-4 orchestration, while OpenReviewer relies on large-scale fine-tuning, making direct benchmarking less meaningful. Thus, we benchmark against SEA for a fair and reproducible comparison.
The integration of graph-based techniques into Retrieval Augmented Generation (RAG) (Lewis et al., 2020) is an emerging research area that aims to enhance the reasoning and factual grounding of LLMs. A recent comprehensive survey by Zhu et al. (2025) provides a foundational taxonomy for this field, categorizing systems based on their database construction, retrieval algorithms, and processing pipelines. Within this framework, AutoRev is a domain-specific Graph-RAG application that generates a hierarchical graph directly from a scientific paper’s text. It employs a learning-based Graph Attention Network (GAT) (Veličković et al., 2018) for retrieval that reduces the context length of input to LLMs, reducing the computational costs.
3. Dataset
@@
The dataset used in our study is borrowed from Yu et al. (2024), comprising papers submitted to various conferences along with multiple peer reviews per paper. Among these, the ICLR 2024 and NeurIPS 2023 datasets are publicly released, though the GPT-4–consolidated reviews for NeurIPS 2023 are not released. Hence, our analysis focuses only on the ICLR 2024 papers. As reported in Yu et al. (2024), the dataset statistics for the ICLR 2024 papers are summarized in Table 1.
Table 1. Dataset statistics for the ICLR 2024 dataset, as reported in Yu et al. (2024).
@@
From the dataset, we utilize only the Summary, Strengths, Weaknesses, and Questions sections from each review, along with the full content of the corresponding paper. Other sections, such as Ratings are excluded, as our framework is designed to provide constructive feedback rather than produce a final decision. The selected sections offer qualitative assessments of the paper’s content and provide valuable insights to guide both authors and reviewers. For our finetuning process, we utilize the GPT-4 consolidated reviews dataset introduced by Yu et al. (2024). This dataset is constructed by using GPT-4 to consolidate multiple human-written peer reviews from OpenReview [^3] https://openreview.net/ into a single, unified review for each paper.
4. Methodology
Figure 2. The system begins with a parsed document from which DPR extracts passages for GNN training. The GNN learns to identify key passages relevant to review generation. The trained GNN selects important passages without requiring human-generated reviews. These passages, combined with a structured prompt, are input to fine-tune an LLM. The fine-tuned LLM generates the final review efficiently, without external grounding.:
In this section, we describe the implementation of our proposed framework. Section 4.1 describes the initial passage extraction using Dense Passage Retrieval (DPR) (Karpukhin et al., 2020). The extracted passages are used as training data in Section 4.3, to train a Graph Neural Network (GNN) that learns to extract relevant passages from a research paper. The trained GNN enables our model to perform competitively on the test set without the need for grounding on human-generated reviews, as discussed in Section 4.4. Post that, we describe our process of fine-tuning LLMs to generate the final reviews, in Section 4.5. An overview of our methodology can be found in Figure 2.
4.1. DPR-based Passage Extraction
@@
This final set of passages represents the most relevant excerpts from the full paper that are semantically linked to the review content. These extracted passages serve as the ground truth labels for training the GNN, as described in Section 4.3
4.2. Graph Construction for Representing a Paper
The idea of graph construction employed in our work is inspired by Zhao et al. (2024). We borrow their idea of representing a document as a hierarchical graph but extend it beyond sentence level. We propose to build a deeper graph, which contributes towards learning more intricate inter-dependencies of passages and sentences in a paper.
Specifically, we formalize the document structure as an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where nodes $\mathcal{V}$ represent text segments with semantically encoded features, and $\mathcal{E}$ denote the hierarchical and sequential relationships that form the document’s structure. Our graph construction is designed to be generalizable across diverse academic paper structures, accommodating arbitrary levels of nesting and heterogeneous section organization.
4.2.1. Node Definitions
@@
  • Heading Nodes ($\mathcal{H}=\{h_{1},h_{2},\dots,h_{m}\}$): Nodes corresponding to top-level section headings.
  • Sub-heading Nodes ($\mathcal{S}_{i}=\{s_{i1},s_{i2},\dots,s_{ik}\}$): Nodes representing sub-headings under each heading $h_{i}$. This structure naturally extends to accommodate multiple levels of nesting, making it suitable for documents with deeper hierarchical organization.
@@
  • Sentence Nodes ($\mathcal{T}_{jk}=\{t_{jk1},t_{jk2},\dots,t_{jkp}\}$): Nodes for individual sentences within each passage $o_{jk}$, tokenized using NLTK (Bird and Loper, 2004).
4.2.2. Edge Definitions
Edges capture both hierarchical and sequential relationships within the document.
  • Root Connections: $(v_{\text{paper}},h_{i})\in\mathcal{E}_{\text{hier}}$ for all $h_{i}\in\mathcal{H}$.
  • Heading-SubHeading Connections:
    $(h_{i},s_{ij})\in\mathcal{E}_{\text{hier}}$ for all $s_{ij}\in\mathcal{S}_{i}$.
@@ 4.2.2. Edge Definitions
    $(s_{ij},o_{jk})\in\mathcal{E}_{\text{hier}}\text{ if }\mathcal{S}_{i}\neq\emptyset;(h_{i},o_{jk})\in\mathcal{E}_{\text{hier}}\text{ otherwise.}$
  • Passage-Sentence Connections:
    $(o_{jk},t_{jkq})\in\mathcal{E}_{\text{hier}}$ for all $t_{jkq}\in\mathcal{T}_{jk}$.
  • Sequential Edges:
    For passages: $(o_{jk},o_{j(k+1)})\in\mathcal{E}_{\text{seq}}$ for $1\leq k<n$.
    For sentences: $(t_{jkq},t_{jk(q+1)})\in\mathcal{E}_{\text{seq}}$ for $1\leq q<p$.
@@
(1) $\mathcal{E}=\mathcal{E}_{\text{hier}}\cup\mathcal{E}_{\text{seq}},$
@@
4.3. GNN Training for Multi Modal Passage Retrieval
Figure 3. Excerpt of a test sample where passages highlighted in cyan colour refer to relevant passages identified by GNN only and yellow colour refer to relevant passages identified by both GNN and DPR:
To train a GNN on the graph constructed in Section 4.2, we take the data generated in Section 4.1, which maps each paper $p_{i}$ to a set of important passages, $\mathcal{O}_{i}=\{o_{1},o_{2},\dots,o_{n}\}$. Each passage is important for generating a review for the paper it is associated with. This mapping is used as a training set for training the GNN. The node embeddings for sentence nodes are initialized with embeddings from a Sentence-Transformer (Reimers and Gurevych, 2019) model, namely all-mpnet-base-v2. The node embeddings for all other nodes are initialized with zero embeddings of dimension $768$, matching that of the Sentence-Transformer model.
Now, a Graph Attention Network (Veličković et al., 2018) is used to train the graph to predict the important passages. After passing the initial representation of the graph nodes through three layers of GATConv (PyTorch Geometric’s implementation of Graph Attention Networks) [^6] https://pytorch-geometric.readthedocs.io/en/latest/generated/torch_geometric.nn.conv.GATConv.html, we apply a non-linear classifier to the final passage node representations. The classification is defined as follows:
  1. (1)
@@ 4.3. GNN Training for Multi Modal Passage Retrieval
We hypothesize that our choice of graph construction which models the semantic representation of nodes and flow of information through structure, allows our model to capture the intricate inter-dependencies of sentences and passages across different sections within an academic paper, and effectively make a choice about which passage(s) is/are important. The dependencies, according to our hypothesis, are learned from the sequential and hierarchical connections that we introduce in our graph construction method. The passing of information from one node to another is enhanced by the GATConv (Equations 2, 3, and 4) layers, which involves the concept of attention to selectively propagate messages in the Message Passing layer of a GNN.
(2) $\mathbf{h}_{i}^{\prime}=\sigma\left(\sum_{j\in\mathcal{N}(i)}\alpha_{ij}\mathbf{W}\mathbf{h}_{j}\right)$
(3) $e_{ij}=\text{LeakyReLU}\Bigl(\mathbf{a}^{T}\bigl[\mathbf{W}\mathbf{h}_{i}\parallel\mathbf{W}\mathbf{h}_{j}\bigr]\Bigr)$
(4) $\alpha_{ij}=\frac{\exp\bigl(e_{ij}\bigr)}{\sum_{k\in\mathcal{N}(i)}\exp\bigl(e_{ik}\bigr)}$
@@
As illustrated in the example from our test set (Figure 3), the DPR-based approach fails to retrieve a key passage related to the main contributions of the paper, whereas our GNN successfully identifies it. Therefore, our multi-modal graph, which leverages both sequential and hierarchical node connections, more effectively identifies the salient passages required to enhance the review generation process. The GNN extracted passages, along with a structured prompt, serve as the input in the LLM fine-tuning stage, as discussed in Section 4.5.
4.4. Grounding Requirements and Inference Efficiency
For practical deployment, our GNN-based method offers a crucial advantage by operating autonomously at inference time. It identifies important passages directly from the paper’s graph structure, eliminating the need for grounding from human-generated reviews which act as queries in the DPR-based approaches, but may be unavailable in real-world scenarios.
Second, by selecting only the passages deemed most relevant, we significantly reduce the input length and associated computational costs for downstream LLM-based review generation. This not only leads to higher performance of LLMs, as outlined by Hsieh et al. (2024) but also leads to higher scores of the generated reviews compared to our baselines (see Section 5). The statistics related to the decrease in number of passages and token length post-extraction by the GNN model are reported in Table 2.
Table 2. Average token length (Avg. Token Len) and Average passage count (Avg. Passages) for a full paper, our best-configuration: GNN $(5,5)$, and our fastest-configuration: GNN $(5,3)$.